OpenStudy (anonymous):

Which shows this equation in standard form? 7 – 3(x – y) = 5x + 2 The choices are - 3x – 8y = 5 8x + 3y = 5 8x – 3y = 5 8x – 3y = –5 I have no idea how to solve this. Could you please answer it and explain to me how you got the answer? All help is very appreciated :)

OpenStudy (anonymous):

7 - 3(x - y) = 5x + 2 -- distribute through the parenthesis 7 - 3x + 3y = 5x + 2 -- subtract 7 from both sides -3x + 3y = 5x + 2 - 7 -- combine like terms -3x + 3y = 5x -5 -- subtract 5x from both sides -3x - 5x + 3y = -5 -- combine like terms -8x + 3y = -5 -- multiply by -1 to make x positive 8x - 3y = 5 <== standard form

OpenStudy (anonymous):

Thank you SO MUCH! :D

OpenStudy (anonymous):

glad to help...do you have any questions ?